Shift summary for day shift, 17/Oct/99

AG (ag@fnal.gov)
Sun, 17 Oct 1999 16:38:17 -0400 (EDT)

Shift Summary for day shift, 17/Oct/99

On Shift: SET, AG

Hours of beam available at all intensities: 8.3
Hours of beam available at nominal intensity: 8.3
Hours of E799 data (special runs not included): 7.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

o stop/start run due to L3 accept problem.

o reboot ktev1.

o change tapes.

========================================================================

Runs ended this shift:
Run 14854 - 3.8M - BEAM_799
Run 14855 - 11.2M - BEAM_799

Tapes brought to FCC this shift:
(including tapes for on-line split)

========================================================================

Accesses during shift:

none.

========================================================================

Overview of shift:

E799 data-taking. Had L3 accept problem once (responded by starting new
run forcing L3 ). Ktev1 gateway died and ktev1 had to be rebooted.
/ktev1e, /ktev2e and /ktev4e disks (ONSPLIT) are 95,96m and 95% full at
the end of the shift, Sydney is working on them.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14854_______0700____Day shift begins.

______________0734____L3 accept alarm. Paged SashaL. Started new run
_______________________forcing L3 executables (no fresh start).

_______36___0827_____Serious alarm -> all scalers are out of limit.
_______________________Goes away by itself during the next spill.
_______________________This alarm goes on from time to time yesterday and
_______________________today - usually goes away next spill.

_______48___0843_____Mounted UPE006 on tps3d2 @ ktev4.

______157___1108_____Beam aborted. Energy conservation in effect.

______163___1117_____Beam is back.

___220-225____________Energy conservation.

_____________1300_____Disk /ktev3a fills up. BobT walks in.

_______308____________Disk /ktev3e fills up, and all "e-disks" close to
_______________________90% as well. ONSPLIT files are causing this. Page
_______________________Sydney. JasonH walks in.

______________1315____Stop run 13855 because ktev1 gateway died. Run
_______________________doesn't stop normally, so we do l3#save and
_______________________drt_fresh_start. We don't know about EOF marks, so
_______________________change tapes (they were ~60% full). Fresh start
_______________________fails because ktev1 is down. Reboot ktev1 (with
_______________________Sydney giving instructions by phone), do fresh
_______________________start.

______________1350____BobT deletes CsI data files for runs before last
_______________________ecal (runs 14832-14845). Usage now is down to 81%.

_______________________Sydney (over the phone) and JasonH are working on
_______________________the filled ONSPLIT disks.

_______________________UPK012 is mounted on tps3d2 @ ktev2.

14856_______1355____Start new run.

______________1425____/ktev1e, /ktev4e are 91 % full. /ktev2e - 92%.
_______________________Sydney is going to be on evening shift, and will
_______________________take care of this.

______________1450____Mounted UPM020 on tps3d2 @ ktev4.

________71___1429____/ktev4e - 95%, /ktev2e - 96%, /ktev1e - 95% ...

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift